A long copper wire with radius of 1.0 mm carries a uniform surface charge density of 5.0 x 10^-6 C/m². 


a. Find the total charge in a 1.0-meter-long section of the wire. 


b.. Find the magnitude of the electric field at a distance of 15 cm from the wire. 



1
Expert's answer
2021-05-31T08:45:29-0400

a. Find the total charge in a 1.0-meter-long section of the wire. First, determine the area:


"A=\pi rL."

The total charge:


"Q=\sigma A=\sigma \pi r L=1.57·10^{-8}\text{ C}."

b. Since the length is infinite, and the radius is 150 smaller than the distance, the field is


"B=\frac{\mu_0 I}{2\pi R}=\frac{\mu_0 Q/t}{2\pi R}=\frac{2.09·10^{-14}}{t}\text{ T},"

where t - time required for the charge Q to move in the wire.